OpenStudy (anonymous):

If 1/2x +1/2(1/2x + 1/2(1/2x +1/2(1/2x + ... = y, then x = ?

jhonyy9 (jhonyy9):

1/2x +1/4x +1/4x +1/4x +... +1/4x =1/2x +n/4x =(2+n)/4x =y 2+n=4xy x=(2+n)/4y